OpenStudy (anonymous):

what is the simplified version of (y^2+y-6) / (y^2+2y-3) The options are a) y-2/y-1 b) y+2/y-1 c) y+2/y+1 d) y-2 / y+1 I've worked it out over and over and all Ive figured out is that it isn't B help please!!

OpenStudy (anonymous):

Yes dear you are right.. This is B..

OpenStudy (anonymous):

Wait so it is b?? I thought it definitely wasnt b?

OpenStudy (anonymous):

Yes it is..

OpenStudy (anonymous):

\[\frac{y^2 + y-6}{y^2 + 2y - 3} = \frac{ (y+3) \times (y-2) }{ (y+3) \times (y-1)}\] Cancel and solve..

OpenStudy (anonymous):

\(\large \color{blue}{\textbf{Oh sorry, it is A}} \)

OpenStudy (anonymous):

thank you!@ thats what i thought

OpenStudy (anonymous):

I had calculated it in mind so mistaken... Never mind.. Welcome..